Two boxes are stacked one on top of the other. One box is 6 feet 6 inches tall, and the other is 4 feet inches 10 tall. How high is the stack?
Answer by srinivas.g(540)

1 feet = 12 inches
height of first box = 6 feet 6 inches = 6*12 + 6 = 78 inches
height of second box = 4 feet 10 inches = 4*12 + 10
= 48+10
= 58 inches
total height = 78+58 =136 inches
or 11 feet 4 inches